A top gun pilot , practising radar avoidance , is manually flying horizontally at 1300kmph 35 m above level ground. suddenly the plane encounters teraain that slopes gently upward at 4.3 degree - difficult to detect visually. How much times does the pilot have to make a correction to avoid flying into ground ?

Grade:12

1 Answers

Puneet Mittal AskiitiansExpert-IITD
13 years ago

Hi Mohit,

tan(4.3 degrees) = 0.0751903777

tan(4.3 deg) = P/B = 35/B

thus B = 35/(0.0751903777) = 465.48 m

time = Horizontal distance which plane has before it hits the terrain / Speed of plane

      = B / (1300 )

B = 465.48 m = 0.46548 Km

thus time = 0.46548 / 1300

              = 0.000358 hr

              = 0.000358 * 3600 sec

              = 1.288 seconds

thus the plane has 1.288 seconds to make a correction to avoid flying into ground

Please feel free to ask as many questions as u have.

Puneet